OpenStudy (anonymous):

what is the equation of the line that has a slope of 4/5 and a y-interception of -1 ?

OpenStudy (nenadmatematika):

it;s a line that goes through the point (0,-1), so: y+1=4/5(x-0) and y=(4/5)x-1

OpenStudy (anonymous):

\[y=\frac{4}{5}x+(-1), y=\frac{4}{5}x-1\]In slope-intercept form, y=mx+b, where m is the slope and b is the y-intercept.